### v1

# Growing Trees and Amoebas’ Replications

Abstract
An amoeba is a tree with a number assigned to each vertex. We describe a natural process of growing trees from a given amoeba and discuss conditions for such a process to be finite.
1 Introduction
We describe a process of growing trees from an amoeba, which is a tree $H$ with a number assigned to each vertex. In a simplified language, this number $m(v)$ tells how many extra edges we will require around a $v\in V(H)$. Given a tree $T$ we look at a fixed copy $H^{\prime}$ of $H$ within $T$, and look whether every $v\in V(H^{\prime})$ is incident with at least $m(v)$ edges within $E(T)\setminus E(H^{\prime})$. If not, we attach the corresponding number of pendent edges to $v$, increasing the size of $T$. A particularly interesting question is whether such growth process can stop or go on forever. This has relations to several known models:
String and graph generation models
are intensively studied in theoretical computer science, formal language theory, and mathematical biology. There are many models of this sort: cellular automata, rewriting systems, graph grammars etc. [13, 6, 10, 12]. In particular, our model resembles graph relabelling systems with forbidden contexts [12]. Most of graph generation models use labelling of graphs. In contrast, we propose a simple rule of tree growth that governed by the structure of subtrees and does not use labels on vertices and/or edges. The resulting process is interesting from graph-theoretical point of view and possibly might impact the area of graph generation models.
Galton-Watson process
is a foundational concept in the realm of probability theory, specifically in the study of family trees and the dynamics of progeny within populations [7]. Introduced by Francis Galton and Henry Watson [3, 7], this stochastic process serves as a mathematical model to examine the transmission of traits, characteristics, or lineage through successive generations of families [7, 11]. The key idea revolves around the probability distribution of offspring within a family, where individuals independently give rise to a random number of descendants [7]. This model has proven particularly valuable in fields such as genetics and demography [3], providing insights into the likelihood of lineage continuation or extinction [7]. The Galton-Watson process remains an essential tool for understanding the unpredictable nature of family structures and the variability in the number of progeny across generations [11].
In contrast to the probabilistic nature of the Galton-Watson process, our process represents a deterministic analog with a distinct feature—its replication criteria are not dictated by a random probability distribution. Instead, the expansion is orchestrated by the existence of a precisely chosen subgraph within the existing tree, acting as a deterministic trigger for the generation of additional leaves or progeny. This deterministic analog introduces a structured and controlled element to the reproductive dynamics, offering a unique perspective on how specific configurations within a family lineage can influence and determine the proliferation of descendants. This departure from stochastic randomness to a deterministic criterion opens up new avenues for exploring the interplay between structure and growth of trees, providing a nuanced understanding of the factors that contribute to the perpetuation or divergence of lineages.
Avoidable paths
were studied in [9, 2, 1]. This concept was generalized further to the case of arbitrary two-rooted graphs in [8] where the so-called inherent graphs were introduced. Verifying inherence is a difficult task. For example, it is not known whether the path $P_{3}$ with edges $st$ and $tv$ is inherent provided $s$, $t$ are the roots. Interestingly, this problem is related to the cops and robbers game [5]. The paper [8] presents necessary conditions for inherence in terms of a pendant extension process, which extends a two-rooted graph by pendant edges.
The process of growing trees from an amoeba generalizes the pendant extension process. It might be useful for further generalizations of avoidability.
This paper is structured as follows. In the beginning of Section 2 we introduce the main concepts of the paper, such as amoeba, its extension, growth, mortality and confinement. In Section 2.1 we relate mortality and confinement conditions while in Section 2.2 we discuss relations between the symmetries of the underlying graphs and the amoeba mortality. In Section 3 we focus on the mortality conditions for amoebas growing in a simplest way, namely, by attaching pendant edges. We identify degree conditions which are necessary for any such amoeba to be immortal in Section 3.1, while in Section 3.2 we describe the immortal amoebas whose underlying graph is a caterpillar. We conclude the paper by Section 4, in which we propose some conjectures as well as directions for future work.
2 Amoebas
An ordinary amoeba [^1] This name was recently given to a different class of graphs [4]. We think that these two classes can not be confused. $A=(T,m)$ is defined as a pair consisting of a tree $T$, together with the root function $m:V(T)\to\mathbb{N}$ mapping every vertex $v$ of $T$ to a non-negative integer $m(v)$ called multiplicity of $v$.
A vertex from $A$ of positive multiplicity is called a root of $A$. Furthermore define $k(A)=\sum_{v\in V(H)}m(v)$.
If not stated otherwise, the amoebas discussed in this paper will always be ordinary.
An $\ell$-extension of amoeba $(H,m)$ is a graph $G$ obtained from $H$ by applying to each vertex $v$ the following procedure: take $m(v)$ disjoint paths $P_{\ell+1}$ and identify one endpoint of each with $v$. The paths are called extension paths.
A copy of amoeba $(H,m)$ in a graph $G$ is an amoeba $(H^{\prime},m^{\prime})$ where $H^{\prime}\leq G$, and $H\cong H^{\prime}$ such that the functions $m,m^{\prime}$ coincide under the corresponding isomorphism. When no fear of ambiguity, we denote amoeba and its copy by the same notation.
Let $A$ be a copy of an amoeba in a tree $G$. Then a $\ell$-growth $G^{\prime}$ of $A$ in $G$ is a tree containing both $G$ and an $\ell$-extension of $A$ and having the minimal possible number of edges.
If, for a copy $A$ of an amoeba in graph $G$, all growths of $A$ in $G$ coincide with $G$, then the copy is called dead. Otherwise, the copy is called alive.
Note that a copy of an amoeba may have non-isomorphic growths, see example in Fig. 1.
Figure 1: Non-isomorhic 3-growths of the same copy of amoeba. The copy is shadowed in grey, the new edge is dashed
For an amoeba $A=(H,m)$, a $\ell$-growth sequence of $A$ starting at a tree $T$ is a sequence of trees $T_{0},T_{1},\dots$ such that $T_{0}=T$ and $T_{i+1}$ is a growth of a copy of $A$ in $T_{i}$, and $T_{i+1}\neq T_{i}$.
For an amoeba $A=(H,m)$, a $\ell$-generation sequence of $A$ is a maximal $\ell$-growth sequence starting at $H$.
If $\ell$ is clear from the context it will be omitted in notation for extensions, growths, generation sequences etc.
Let $\ell$ be the extension parameter and $A$ be an amoeba. It is immortal if it admits an infinite generation sequence, otherwise it is mortal. Amoeba $A$ is strongly immortal if all its generation sequences are infinite. Amoeba $A$ is partially immortal if it admits an infinite growth sequence.
A tree $T$ is confining for an amoeba $A=(H,m)$ if it admits a copy of $A$ and all copies of $A$ in $T$ are dead.
An amoeba is totally confined if for each tree $T$ containing a copy of $A$ there exists a confining tree $T^{\prime}$ containing a copy of $T$.
Remark 1.
Confining trees can be replaced by confining graphs. Totally confined amoebas can be defined in this more general setting too.
Remark 2.
In [8], general two-rooted graphs are considered in the extension process. Since attaching paths to a graph does not change its set of cycles, this process is finite for non-trees. As we mainly study mortality, we restrict our attention to the case of trees.
2.1 Initial observations
Note that a copy of an amoeba may have non-isomorphic growths, which make things complicated. So, we start from study of this effect to show that it appears for small amoebas (with respect to $\ell$) and in early members of growth sequences. An example in Fig. 1 can not be seen in generation sequences, as we will prove. It is caused by the minimality condition.
Let $T^{\prime}$ be an $\ell$-growth of a alive copy of an amoeba $A=(H,m)$ in a tree $T$, and let $A^{\prime}$ be the corresponding extension of $A$ contained in $T^{\prime}$, and $P$ be a path $v_{0},v_{1},\dots,v_{\ell}$ of the extension starting at a vertex $v_{0}$ such that $\{v_{\ell-1},v_{\ell}\}\notin E(G)$. Take a connected component of the forest $T^{\prime}-H$ containing $v_{1}$. It is a tree $T^{\prime\prime}$ rooted at $v_{1}$, which is called an area of $P$. Due to the minimality condition, the depth of $T^{\prime\prime}$ (the maximal length of a path from the root to a leaf) is $\ell-1$. Moreover, $v_{1},\dots,v_{\ell}$ is a unique path of length $\ell-1$ from the root $v_{1}$. Note that other edges in $T^{\prime}$ are the edges of the original tree $T$. In this way we come to the following conlusion.
Proposition 1.
An intersection of an area of an extension path and the original graph is a tree of depth at most $\ell-1$.
Given a tree $T$ and its subtree $T^{\prime}$, there is a partial order on $V(T)\setminus V(T^{\prime})$ defined as follows: $u\leq^{T^{\prime}}v$ if $u$ belongs to the path from $v$ to $T^{\prime}$. For each $v\in V(T)\setminus V(T^{\prime})$ the set $\{u:v\leq^{T^{\prime}}u\}$ induces a tree ${\mathcal{I}}_{v}^{T^{\prime}}$ rooted at $v$. We omit the mentioning of $T^{\prime}$ in the exponent when it is clear from context.
Lemma 2.
Let $T_{0},T_{1},\dots,T_{k}$ be an $\ell$-growth sequence for an amoeba $A=(H,m)$ and $v\in V(T_{k})\setminus V(T_{0})$ such that ${\mathcal{I}}_{v}^{T_{0}}$ has the depth at most $\ell-1$. If for some vertex $u$ in ${\mathcal{I}}_{v}^{T_{0}}$ there are two different paths $P$ and $P^{\prime}$ of equal lengths from $u$ to a leaf, then
[…]
Proof.
W.l.o.g. we can assume that $P$ and $P^{\prime}$ have no common edges. Indeed, otherwise we can change the vertex $u$ by the last common vertex of $P$ and $P^{\prime}$. Choose the maximal $i$ such that at least one edge of $P\cup P^{\prime}$ is not in $T_{i}$. W.l.o.g. we can assume that $e=(u_{0},u_{1})$ is the last edge of $P$ and $e$ is not in $T_{i}$. It means that $e$ is added in a growth in $T_{i}$ of a copy $A=(H,m)$ of the amoeba. So an extension of $A$ contains an extension path from a root $r$ of $A$ to $u_{1}$. We claim that $u$ is not on the path from $r$ to $u_{1}$ that has no common edges with $P^{\prime}$. Indeed, otherwise $P^{\prime}$ does not intersect $H$ and the extension path can be changed to the path containing $P^{\prime}$, contradicting the minimality condition, see Fig. 2(a). Also, $r$ is not on $P$. Indeed, otherwise the extension should add more edges, since the length of the whole path $P$ is less than $\ell$, see Fig. 2(b).
Refer to caption Refer to caption
(a) (b) $a<\ell$
Refer to caption Refer to caption
(c) (d)
Figure 2: Four ways of extension
In contrast, the next two options are possible and mutually exclusive:
  • vertices $u$, $u_{1}$, $r$ are end-points of paths starting at a vertex $w$, Fig. 2(c);
  • the path from $r$ to $u_{1}$ has a common edge with $P^{\prime}$, see Fig. 2(d).
In both cases $A$ is contained in the subtree rooted at $r$ since the path from $r$ to $u_{1}$ has no common edges with $A$. This subtree is inside ${\mathcal{I}}^{T_{0}}_{u}$ so its depth is at most $\ell-2$, which implies that the diameter of $H$ is at most $42\ell-$ 4. Also, any path starting at $r$ and containing $w$ and $u$ should has the length at most $\ell-1$, by the minimality condition. Since $T_{i}$ contains $T_{0}$ and the path from $r$ to $T_{0}$ has at least 3 edges, the diameter of $T_{0}$ is at most $62\ell-$ 6. ∎
Lemma 2 puts restrictions on extensions using ‘new’ edges only. Ambiguous $\ell$-growths are possible in generation sequences even if $H$ has the diameter at least $62\ell-$ 6. See example in Fig. 3.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 3: An amoeba admitting ambiguous $3$-extensions in a generation sequence
It is open if these three versions of immortality are equivalent.
Figure 4: An immortal amoeba with one root with respect to $2$-extensions. Next step uses the left lower $P_{5}$ for extensions. Repetitions of the process continue forever.
It is clear that the last tree in any generation sequence is confining. So mortality implies the existence of confining tree. We do not know whether the opposite is true. We are able to prove the opposite for the values of the extension parameter $\ell\in\{1,2\}$.
Lemma 3.
For extension parameter $\ell\in\{1,2\}$, an amoeba is mortal if and only if there exists a confining tree for it.
Proof.
In one direction it is obvious.
In other direction: take a confining tree and a copy of the amoeba. Start a generation sequence from this copy. Note that $1$- and $2$-growths of a copy of the amoeba in any tree are isomorphic since in a rooted tree of depth $\leq 1$ all leaves belong to the same orbit of automorphisms of the tree. It implies that at each step of extension new edges can be chosen from the confining tree. Therefore the sequence is finite. ∎
Remark 3.
In general case, strong immortality implies immortality but the converse is open. For the extension parameterr $\ell\in\{1,2\}$ the converse follows from Lemma 3. If an amoeba is not strongly immortal then the last tree of its finite generation sequence is a confining tree for the amoeba. Thus any its generation sequence is finite.
In the proof of the next theorem we will use the notion of full $t$-ary tree of depth $\ell$. To this end we denote by $T^{t,\ell}$ a tree where every node on depth less than $\ell$ has $t$ children while every node on depth $\ell$ is a leaf.
Theorem 4.
An amoeba is not partially immortal if and only if it is totally confined, regardless of its extension parameter.
Proof.
Let $\ell$ be the extension parameter.
In one direction it is easy. Suppose an amoeba is not partially immortal. Take a tree $T$ and any $\ell$-growth sequence $T=T_{0},\dots,T_{k}$ starting at $T$. It is finite and the last tree $T_{k}$ in the sequence confines the amoeba and contains $T$.
For the other direction, take a tree $T_{0}$ and a totally confined amoeba $A=(H,m)$. Let $\delta=\max_{v\in V(H)}\big{(}d(v)+m(v)\big{)}$, $d=\mathop{\mathrm{diam}}(T_{0})$. Choose a confining tree $T$ that contains a full $\delta$-ary tree $T^{\delta,3\ell+2d}$ of sufficiently large depth, say, $3\ell+2d$. Choose a copy of $T_{0}$ containing the center of $T^{\delta,3\ell+2d}$. Start $\ell$-growth sequence $T_{0},T_{1},\dots$ from this particular copy. If $\mathop{\mathrm{diam}}(T_{i})\leq 2\ell-6$ then any copy of $A$ in it is at distance at most $62\ell-$ 6 from the center of $T^{\delta,3\ell+2d}$. Thus any growth of $T_{i}$ can be embedded to $T$. If an $\ell$-growth of $T_{i}$ uses edges of $T_{0}$, then the corresponding copy of $A$ is at distance at most $2d$ from the center of $T^{\delta,3\ell+2d}$. Thus, again, the growth can be embedded to $T$. If an $\ell$-growth of $T_{i}$ does not use edges of $T_{0}$ and $\mathop{\mathrm{diam}}(T_{i})>2\ell-6$, then Lemma 2 implies that all such growths are isomorphic. So we apply the same arguments as in the proof of Lemma 3 and conclude that $T_{i+1}$ can be embedded in $T$. Therefore the sequence is finite. ∎
Conjecture 1.
Let $\ell$ be the extension parameter, let $A=(H,m)$ be an amoeba. If $\Delta(H)>k(A)+2$ then $A$ is totally confined.
2.2 Orbit lemma
Fix an amoeba $A=(H,s_{1},\dots,s_{k})$. For a vertex $v\in V(H)$, we define the orbit of $v$, denoted by $\mathop{\mathrm{Orb}}(v)$, as the set of vertices $w\in V(H)$ such that there exists an automorphism of $H$ mapping $v$ to $w$. The completion $[A]$ of amoeba $A$ is defined as
$[A]=(H,m^{\prime}),\qquad m^{\prime}(v)=\max_{w\in\mathop{\mathrm{Orb}}(v)}m(w).$
We call a pair of amoebas $A=(H,m)$ and $A^{\prime}=(H,m^{\prime})$ equivalent (to each other) if $[A]=[A^{\prime}]$. Using this notation one can easily obtain the following claim.
Lemma 5.
Let $\ell$ be the extension parameter and $A=(H,m)$ be an amoeba. If $\ell\in\{1,2\}$ then $A$ is immortal if and only if its completion is.
For other values of $\ell$, an amoeba is partially immortal if and only if its completion is.
Proof.
We start from the case $\ell\in\{1,2\}$. Suppose that the completion $A^{\prime}$ of $A$ is mortal, that is, it admits a confining tree $T^{\prime}$. It remains to observe that $T^{\prime}$ confines also $A$, so $A$ is mortal by Lemma 3.
Now suppose by contradiction that $A$ is mortal and $A^{\prime}=(H,m^{\prime})$ is not. Lemma 3 implies that there exists a confining tree $T$ of $A$. Since $T$ does not confine $A^{\prime}$ there exists a simplicial copy of $A^{\prime}$ in $T$, that is, for some vertex $v\in V(H)$ it holds
$d_{T}(v)-d_{H}(v)\leq m^{\prime}(v).$
Since $m^{\prime}(v)=\max_{w\in\mathop{\mathrm{Orb}}(v)}m(w)$, there is a vertex $w\in\mathop{\mathrm{Orb}}(v)$ such that $m(w)=m^{\prime}(v)$. Since $v$ and $w$ are in the same orbit of $H$, there exists a copy of $A$ such that
$d_{T}(v)-d_{H}(v)\leq m(v),$
which contradicts the choice of $T$.
For $\ell>2$, similar arguments show that an amoeba is partially immortal if and only if its completion is. The only difference is to use Theorem 4 instead of Lemma 3. ∎
3 Case of $1$-extensions
Understanding the mortality of general amoebas seems to be very difficult. In this section we focus on the mortality conditions for amoebas which grow in the simplest of ways – by attaching pendant edges. In our notation such amoeba growth corresponds to the case of $1$-extensions. In [8] authors show that for the case $k(A)=2$, any immortal amoeba is necessarily subcubic. The following subsection generalizes those results to the case of arbitrary $k(A)$.
3.1 Degree conditions
For the rest of this subsection fix an immortal amoeba $A=(H,m)$. Let $(d(v),m(v))$, $v\in H$, be the sequence of vertex degrees and multiplicities of the roots in $H$. We also define
$\tilde{d}(v)=d(v)+m(v)\qquad\text{and}\qquad m=\max_{v\in V(H)}\tilde{d}(v).$
Fix a $1$-generation sequence $G_{0},G_{1},\dots$. The immediate corollary of minimality of $\ell$-generation is the following claim.
Claim 6.
If a vertex $v$ of $G_{i}$ is a root of a copy of the amoeba to be extended and $d_{G_{i+1}}(v)>d_{G_{i}}(v)$ then $d_{G_{i}}(v)<\tilde{d}(v)$.
Proof.
If $d_{G_{i}}(v)\geq\tilde{d}(v)$ then an extension uses existing edges incident to $v$ due to minimality condition. ∎
Corollary 7.
$d_{G_{i}}(v)\leq m$ for all $i$ and $v\in V(G_{i})$.
Corollary 8.
For all $i$ and sufficiently large $j$, vertices of a copy extending at step $j$ do not intersect $V(G_{i})$.
These facts put restrictions on degrees of vertices in $H$. Recall that we add pendant edges. So, initially a degree of a new vertex $v$ is 1. Then it can be increased by extension of a copy having a root at $v$. Thus $\min_{v:m(v)>0}d(v)=1$.
To catch the process of degree increasing, we introduce an auxiliary directed graph $D$. The vertices of $D$ are positive integers up to $m$. A pair $(x,y)$ is an edge in $D$ if $y=d(v)+m(v)$ and $d(v)+m(v)>x\geq d(v)$ for some $v\in V(H)$.
Let $q$ be the maximal integer reachable from 1 in $D$. By definition of $q$, the inequality $q\leq m$ holds. For an immortal amoeba $m=d(v^{*})+m(v^{*})$ should be reachable because the degree $d(v^{*})$ should appear in some new vertex (i.e. a vertex from $V(G_{i})\setminus V(G_{0})$) and all copies containing this vertex at some moment have enough space for the extension. This implies the following theorem.
Theorem 9.
Let $A=(H,m)$ be an immortal amoeba. Then
  1. 1.
    $q=m$;
  2. 2.
    $d_{H}(v)\leq m$ for all $v\in V(H)$.
Claim 10.
$m\leq 1+k$.
Proof.
Let $x_{1},x_{2},\dots,x_{t}$ be a directed path in $D$. By definition of $D$, $x_{i}=d(v_{i})+m(v_{i})$ and $x_{i-1}\geq d(v_{i})$, for $i>1$. Then
$x_{1}\leq 1;$
$x_{2}=d(v_{2})+m(v_{2})\leq 1+m(v_{2});$
$\dots$
$x_{t}\leq 1+\sum_{j=2}^{t}m(v_{j})\leq 1+\sum_{v\in V(H)}m(v)=1+k.$
The last inequality follows from the fact that each $x_{j}$ can appear along the path only once. ∎
Corollary 11.
Let $A=(H,m)$ be amoeba. If $\Delta(H)>1+k$, then $A$ is mortal.
Corollary 12.
Let $A=(H,m)$ be an immortal amoeba such that $k=1$. Then $\Delta(H)\leq 2$.
3.2 Mortality criteria for caterpillars
In this subsection a $1$-extension of a given amoeba will be referred to simply as extension. Furthermore, throughout this section we assume that the multiplicity function $m$ takes values 00 and $1$ only. A caterpillar $C(d_{1},d_{2},\dots,d_{\ell})$ is a tree constructed as follows. We start with a central path $P_{\ell}=(p_{1},\dots,p_{\ell})$, and afterwards, for each $i\in\{1,\dots,\ell\}$, we add $d_{i}$ pendant vertices to $p_{i}$. In addition, we require $d_{1}=d_{\ell}=0$. A caterpillar is said to be symmetric if $d_{i}=d_{\ell-i+1}$, for all $i$. The family of amoebas $(H,m)$ where $H$ is a caterpillar and $m$ respects the above mentioned properties is denoted by $\mathcal{A}$. In this section we characterize immortal amoebas from $\mathcal{A}$.
Observation 13.
Let $H$ be a caterpillar with central path $P\subseteq V(H)$, and let $A=(H,m)$ be an amoeba such that $\operatorname{supp}(m)\subseteq P$. An extension of $A$ gives rise to a caterpillar.
Figure 5: A caterpillar $(C(0,2,2,3,0),m)$ and its extension $C(0,0,2,3,4,0)$.
Proof.
It is enough to observe that the central path of the obtained graph corresponds to $P$, possibly with additional vertex on either endpoint of $P$ (also see fig. 5). ∎
We now give necessary and sufficient condition for an amoeba from $\mathcal{A}$ to be shifted ”to the right”. But first we need some definitions.
Definition 1 (of slow integer sequences).
Let $S=(d_{1},d_{2},\dots)$ be an integer sequence. We say that $S$ is
  • slowly decreasing, if $d_{i}-d_{i-1}\geq-1$ for all $d_{i},d_{i-1}$;
  • slowly increasing, if $d_{i}-d_{i-1}\leq 1$ for all $d_{i},d_{i-1}$;
  • slow if it is either slowly increasing or slowly decreasing.
Definition 2 (of slow amoebas).
For an integer sequence $S$ let $(p_{1},\dots,p_{\ell})$ be the central path of $C(S)$ and $(C(S),m)$ be an amoeba from $\mathcal{A}$. The amoeba $(C(S),m)$ is said to be
  • slowly decreasing if $S$ is slowly decreasing and the roots are placed at $p_{i}$ whenever $d_{i}-d_{i-1}=-1$, and also in $p_{\ell}$;
  • slowly increasing if $S$ is slowly decreasing and the roots are placed at $p_{i-1}$ whenever $d_{i}-d_{i-1}=1$, and also in $p_{1}$;
  • slow if it is either slowly increasing or slowly decreasing.
Theorem 14.
An amoeba $A\in\mathcal{A}$ is immortal if and only if $[A]$ is slow.
Proof.
Let $S=(d_{1},d_{2},\dots)$ be the degree sequence of the central path of $[A]$. To prove the ‘if direction’ assume wlog that $[A]$ is slowly decreasing. We define a sequence of graphs $\mathcal{G}=(G_{1},G_{2},\dots)$, starting with $G_{1}=C(S)$. (Informally speaking, $\mathcal{G}$ is a sequence of ”right shifts” of $[A]$.) Every member of this sequence is a caterpillar graph, hence, the vertex notation (vertices $p_{1},\dots$) is well defined.
Within $G_{i}$ define a set of vertices $\mathcal{C}_{i}\subseteq V(G)$ as $\mathcal{C}_{i}=\{p_{i},\dots,p_{i+|S|}\}$.
If $G_{i}$ admits a copy $H$ of $C(S)$ containing $\mathcal{C}_{i}$ then $G_{i+1}$ is obtained from $G_{i}$ by extending $H$.
Due to Lemma 3 of the ‘if direction’ of the statement it is enough to prove that $\mathcal{G}$ is infinite. Assume for contradiction that $\mathcal{G}$ has (a finite) length $k$. By definition $G_{1}=C(S)$ contains $\mathcal{C}_{1}$, hence, $k\geq 2$.
Furthermore, $G_{k}$ does not admit a copy of $C(S)$ containing $\mathcal{C}_{k}$. In other words, either
  1. (i)
    there is an $i$ such that $d_{G_{k}}(p_{k+i})<d_{i}+2$ with $i<|S|$, or
  2. (ii)
    $d_{G_{k}}(p_{k+i})<d_{i}+1$ with $i=|S|$.
Note that (ii) implies $d_{G_{k}}(p_{k+|S|})=0$ which is impossible. Let us fix an $i$ satisfying (i).
By construction, there exists a copy $H$ of $C(S)$ in $G_{k-1}$ containing $\mathcal{C}_{k-1}$ whose extension gives rise to $G_{k}$. By (i) we have $d_{i}-d_{i-1}<0$, furthermore $d_{i}-d_{i-1}=-1$ since $S$ is slowly decreasing. However, due to the extension of $[A]=(H,m^{\prime})$ in $G_{k-1}$, we are guaranteed that $m^{\prime}(p_{k+i})>0$, and hence $d_{G_{k}}(p_{k+i})-2\geq d_{i}$, which is a contradiction.
To prove the ‘only if direction’ of the claim assume that $[A]=(H,m^{\prime})$ is neither slowly increasing nor slowly decreasing. This can be due to one of the following reasons. There exist:
  1. 1.
    $i$ and $j$ such that $d_{i}-d_{i-1}<-1$ while $d_{j}-d_{j-1}>1$;
  2. 2.
    $i$ and $j$ such that $d_{i}-d_{i-1}<-1$, and $d_{j}-d_{j-1}=1$ while $p_{j-1}$ is not a root;
  3. 3.
    $i$ and $j$ such that $d_{i}-d_{i-1}=-1$, and $d_{j}-d_{j-1}>1$ while $p_{i}$ is not a root;
  4. 4.
    $i$ and $j$ such that $d_{i}-d_{i-1}=-1$ and $d_{j}-d_{j-1}=1$, while neither $p_{i}$ nor $p_{j-1}$ are roots.
Consider the central path $(p_{1},\dots,p_{\ell})$ in an arbitrary copy of $[A]$ in $G_{1}$. In all four above cases the following claim holds:
Lemma 15.
For all copies of $[A]$ in $G_{2}$, the sets $\{p_{2},\dots,p_{\ell-1}\}$ coincide.
Proof.
Indeed, any path of length $\ell-1$ in $G_{2}$ contains a pendant vertex either from $V(G_{2})\setminus V(G_{1})$ or from $G_{1}$, and in the latter case it is adjacent to either $p_{2}$ or $p_{\ell-1}$. In fact, the former case of this alternative is possible only if the amoeba is slow. Otherwise, one of the cases listed above holds. ∎
Denote by $(p^{\prime}_{1},\dots,p^{\prime}_{\ell})$ the vertices of $G_{2}$ which correspond to $(p_{1},\dots,p_{\ell})$ of $A$. By Lemma 15, for $1<i\leq\ell$, we have that either
$p^{\prime}_{i}=p_{i-1},\text{ or}\quad p^{\prime}_{i}=p_{i},\text{ or}\quad p^{\prime}_{i}=p_{\ell-i+1},\text{ or}\quad p^{\prime}_{i}=p_{\ell-i}.$
In all four cases, we obtain a contradiction with the vertex degrees’ conditions in $G_{2}$.
To finish the proof, take a copy $A^{\prime}$ of $[A]$ in $G_{2}$ such that its central path contains a pendant vertex $v$ from $G_{1}$ adjacent to $p_{\ell-1}$; the case when it is adjacent to $p_{2}$ is proved similarly. To have enough room for $A^{\prime}$ to grow, we add one pendant edge to $v$ only. Notice that no new copies of $[A]$ appear, thus, $[A]$ is mortal, and due to Lemma 5, so is $A$. ∎
4 Directions for future work
As announced in Section 2, we conjecture that the three notions of immortality are equivalent.
Conjecture 2.
The following properties of an amoeba $A$ are equivalent:
  • $A$ is immortal.
  • $A$ is strongly immortal.
  • $A$ is partially immortal.
Deciding mortality looks hard (undecidable); perhaps one can model a general Turing machine by amoebas. For this purpose a more general notion of an amoeba colony might be useful. An amoeba colony is a set of ordinary amoebas $\mathcal{C}=\{A_{1},\dots,A_{n}\}$. An $\ell$-growth $G^{\prime}$ of $\mathcal{C}$ in $G$ is an $\ell$-growth of some $A_{i}\in\mathcal{C}$. The remaining definitions are the same as for the case of ordinary amoebas. So, in construction of $\ell$-growth sequence for an amoeba colony, one can choose an alive amoeba to be extended. It gives more freedom and we conjecture that such a process can simulate a universal computation.
Conjecture 3.
The following algorithmic problem is undecidable.
Input: an integer $\ell$, a tree $T$ and an amoeba colony $\mathcal{C}$.
Question: is there an infinite growth sequence for $\mathcal{C}$ starting from $T$?
The following generalizations might be of independent interest.
A more general type of an extension of an amoeba is to attach to each vertex a specified tree. A general-type amoeba is $(T,\pi)$, where $\pi$ is a root function from $V(T)$ to rooted trees (the empty tree is allowed). All other definitions are naturally extended to the case of general-type amoebas.
A general-type amoeba is called absolutely mortal if it is mortal for any root function $\pi$. Characterization of absolutely mortal amoebas is an interesting open question in this area.
Another way of generalization is to substitute trees by arbitrary graphs, possibly disconnected. We know a little about this general problem. But it also seems to be interesting.
As already mentioned in Remark 1, we can extend the definition of confining trees to general graphs. This would naturally generalize the concept of totally confined amoebas. From our previous results [8] it follows that there exist amoebas which do not admit confining trees but can be confined by a general graph. For various amoebas, the confining graphs are cages: Petersen graph, Heawood graph, McGee graph, Tutte–Coxeter graph. The confinement of general amoebas might lead to interesting results about cages and other graphs having rich structure.
Acknowledgements.
Authors are grateful to Prof. Martin Milanič for valuable discussions and observations.
This research was partially prepared within the framework of the HSE University Basic Research Program. This work is supported in part by the Slovenian Research Agency (research program P1-0383 and research projects N1-0160, N1-0209, J1-3003, J1-4008 and J5-4596).